OpenStudy (anonymous):

A savings account compounds interest, at a rate of 17%, once a year. John puts $1,000 in the account as the principal. How can John set up a function to track the amount of money he has

OpenStudy (anonymous):

A)A(x) = 1000(17)x where 17 is the interest rate B)A(x) = 1000(.17)x where .17 is the interest rate C) A(x) = 1000(1 + .17)x where .17 is the interest rate D)A(x) = 1000(1 + 17)x where 17 is the interest rate